Give that the variance of a set of scores is 100 and the coefficient of variation is 20%, determine the mean for the scores
a.50
b.500
c.20
d.200
1
Expert's answer
2020-11-24T16:54:07-0500
"coefficient\ of\ variation=\dfrac{\sigma}{\bar{x}}"

Then


"\bar{x}=\dfrac{\sigma}{coefficient\ of\ variation}="


"=\dfrac{\sqrt{100}}{0.2}=50"

a.50